OpenStudy (anonymous):

how do i simplify these expressions? -20 - (-5)*(2^2) (1/4)^3

OpenStudy (anonymous):

Exponent first, evaluate the 2^2 next, then take put together with the fraction to get - ((-5x4)/(4^3))-20 then \[-\frac{ -5*4 }{ 64 }-20\] multiply -5 and 4 \[-\frac{ -20 }{ 64 }-20\] simplify \[\frac{ 5 }{ 16 }-20\] to subtract 20 you need to get the same denominator by finding a number that both 20 and 16 go into then add them together and simplify. If you still can't figure it out let me know.

OpenStudy (anonymous):

okay thank you!